OpenStudy (anonymous):

After a 70% reduction, you purchase a new soft drink machine on sale for $168. What was the original price of the soft drink machine?

OpenStudy (anonymous):

168/.70 = 240 check 240 * .70 = 168 ok

hero (hero):

\[\% \Delta = \frac{|a - b|}{a} \times 100\] a = original price b = sales price

hero (hero):

\[70 = \frac{|a - 168|}{a} \times 100\]

OpenStudy (anonymous):

I some how just got 560

OpenStudy (anonymous):

but 70% of 560 is 392 so thats not right

OpenStudy (anonymous):

its not

hero (hero):

There's another formula you can use

hero (hero):

x - .70x = y x = original price y = sales price

OpenStudy (anonymous):

560 is correctt

hero (hero):

x(1-.7) = 168 .3x = 168 x = 168/.3 x = 560 Sorry, sweetie (it really is 560)

hero (hero):

I just proved it using two formulae

hero (hero):

@JuanitaM

OpenStudy (anonymous):

agreed

hero (hero):

How come I didn't receive any medals?

hero (hero):

thanks

OpenStudy (anonymous):

I gave medals